Grade 12Mechanics

A rope if length L and mass M is hanging from a rigid support. The tension in the rope is at a distance x from rigid support is :
(a)Mg
(b)(L-x/L)
(c)L/L-x)Mg
(d)(X/L)Mg

dushyanta kashyap
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
if the chain is held vertically downward two forces will be acting along the surface of chain
one of them is the mass of the chain that is mg acting vertically downward and other is the tension of chain acting vertically upward and canceling the effect of mass pulling it downward
since acceleration is zero so 
F(net)=0
so mass of chain is m=(L-x)
T=mg
m=(L-x)/L*m
T=(L-x)/LM*g